Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn, all entries  Not logged in ELOG logo
ID Date Author Group Subject Text Attachments
  69   14 May 2025, 09:39 Jan GruenertPRCStatusDelayed start of SA1 (and SA3) beam delivery

The
beam delivery to the North branch (SASE1
/ SASE3) was delayed by 24h compared to the
 JG_2025-05-14_um_09.37.38.png 
  68   13 May 2025, 07:00 Raphael de WijnSPBExperiment summaryEnd of in-house beamtime

X-ray delivery


9.3 keV, 2.2-2.5 mJ pulse
  
  67   09 May 2025, 06:59 Raphael de WijnSPBShift summary
Shift
summary in-house night 2

X-ray delivery
  
  66   08 May 2025, 04:56 Raphael de WijnSPBShift summaryShift summary in-house beamtime night 1

X-ray delivery


9.3 keV, 1.7-2.2 mJ pulse
  
  65   05 May 2025, 22:41 Tokushi SatoSPBShift summaryX-ray
delivery


8.0 keV, 1300-1600 uJ pulse
  
  64   04 May 2025, 23:14 Tokushi SatoSPBShift summaryX-ray
delivery


8.0 keV, 1300 uJ pulse
  
  63   03 May 2025, 23:13 Tokushi SatoSPBShift summaryX-ray
delivery


8.0 keV, 1600 uJ pulse
  
  62   03 May 2025, 06:47 Tokushi SatoSPBShift summary**X-ray delivery**
- 8.3 , 1600 uJ pulse energy

**Optical laser delivery**
  
  61   02 May 2025, 08:07 Andreas KochXPDIssueConcerning eLog entry #59, SPB PopIn:

The problem of the motor position,
  
  60   01 May 2025, 23:35 Tokushi SatoSPBShift summaryX-ray
delivery


8.3 / 9.3 keV, 1400 uJ
  
  59   01 May 2025, 07:25 Romain LetrunSPBIssueSPB pop-in monitor was not dricing to the
switch when we retracted it to the "out
" position with the MDL this morning.
 Screenshot_from_2025-05-01_07-21-57.png 
  58   30 Apr 2025, 22:46 Romain LetrunSPBShift summaryX-ray delivery


9.3 keV, 1800 uJ pulse
energy

1 - 202 pulses at
  
  57   26 Apr 2025, 06:55 Tokushi SatoSPBShift summaryStaff: PV, TS, JK + external collaborators 

X-ray delivery:

10 keV, ~ 800 uJ, 128 pulses, tuning
  
  56   25 Apr 2025, 06:48 Tokushi SatoSPBShift summaryX-ray delivery:

10 keV, ~ 850 uJ, 128 pulses, tuning
was needed for 1 hour 
  
  55   24 Apr 2025, 08:42 Konopkova ZuzanaHEDIssueself-seeding level in SA2 is currently
very low (~100 uJ, about 50% in the pedestal),
need more information on how to proceed -
  
  54   24 Apr 2025, 06:56 Jayanath KoliyaduSPBShift summaryX-ray delivery:

10 keV, ~ 900 uJ, 128 pulses, tuning
was needed for few hours
  
  53   17 Apr 2025, 01:39 Chan KimSPBShift summaryX-ray delivery


18.8 keV, 260 uJ pulse
energy


Optical
  
  52   14 Apr 2025, 23:08 Tokushi SatoSPBShift summaryX-ray delivery:

20 keV,  250-400 uJ, 250 pulses,
stable delivery
  
  51   14 Apr 2025, 09:50 Jan GruenertXPDStatusBeam Downtime

During last night, apparently the
photon beam was down in all SASE beamlines
 JG_2025-04-14_um_09.43.03.png 
  50   13 Apr 2025, 22:55 Tokushi SatoSPBShift summaryStaff: PV, TS, JK, FK, + users

X-ray delivery:

20 keV, ~ 300-400 uJ, 250 pulses,
  
  49   12 Apr 2025, 23:15 Tokushi SatoSPBShift summaryX-ray delivery:

20 keV, ~ 500 uJ, 246 pulses, stable
delivery
  
  48   11 Apr 2025, 23:14 Tokushi SatoSPBShift summaryX-ray delivery:

20 keV, ~ 400 uJ, 246 pulses,
stable delivery
  
  47   10 Apr 2025, 23:17 Tokushi SatoSPBShift summaryX-ray delivery:

20 keV, ~ 400 uJ, 128 pulses, stable
delivery 
  
  46   09 Apr 2025, 22:46 Tokushi SatoSPBShift summaryX-ray delivery:

20 keV, ~ 400 uJ, 128 pulses, stable
delivery 
  
  45   09 Apr 2025, 10:22 Jan GruenertPRCSA3 statusSASE3 Apple-X undulator cell 26
issue - SOLVED

During the night, an issue appeared
  
  44   09 Apr 2025, 09:41 Jan GruenertPRCIssueSASE1 XTD2 attenuator arm issue

On monday 7.4.2025, one of the
SASE1 attenuator arms (SA1_XTD2_ATT/PNACT/A9)
  
  43   07 Apr 2025, 15:35 Jan GruenertPRCStability IssueSASE1 pulse energy drops slowly
three times

SASE1
XGM detects three times during the morning
 JG_2025-04-07_um_15.34.41.png 
  42   07 Apr 2025, 07:19 Harald SinnPRCSA1 status7:15 SPB reports large background from
non-lasing SASE3 bunches. BKR is informed.  

7:50 BKR analysed situation. Everything
  
  41   06 Apr 2025, 06:33 Harald SinnPRCSA2 status5:49 HED called and they have problems
with the Laser Sync signal. A change of laser
delay is not possibe. We supsect that it
 LBsync1.jpgLBsync2.jpgLBsync3.jpgLBsync4.jpgLBsync5.jpg 
  40   03 Apr 2025, 07:34 Harald SinnPRCSA1 status7:20 FXE reports a problem with laser synchronisation
(yellow warning light and could not change
delay). BKR informs the DESY laser sync group. 
  
  39   27 Mar 2025, 22:57 Tokushi SatoSPBShift end**X-ray delivery:**
* 16 keV, ~upto 1 mJ, 160 pulses
**Optical Laser:**
  
  38   17 Mar 2025, 06:51 Romain LetrunSPBIssueThe status of the FXE shutter in the Public
status display DOOCS panel is incorrect.
Here appears as opened, when it is actually
 Screenshot_from_2025-03-17_06-45-52.png 
  37   15 Mar 2025, 19:49 Harald SinnPRCStatusMID called and reported issue with Dachs
terminals at the user labs (probably no power).
TS was informed. Because the lab are
  
  36   12 Mar 2025, 23:14 Raphael de WijnSPBShift summary
SPB/SFX
shift summary user experiment 6933 (setup
day)
  
  35   04 Mar 2025, 07:05 Tokushi SatoSPBShift summaryX-ray delivery

    up to 188/352
pulses per train (~1 mJ pulse energy)@13.4
  
  34   03 Mar 2025, 07:05 Tokushi SatoSPBShift summaryX-ray delivery

    up to 188 pulses
per train (~1 mJ pulse energy)@13.4 keV,
  
  33   02 Mar 2025, 06:56 Tokushi SatoSPBShift summaryX-ray delivery

    up to 175 pulses
per train (~1 mJ pulse energy)@13.4 keV,
  
  32   01 Mar 2025, 11:53 Martin DommachPRCIssueDue to issues
SCS has with pulses from SASE1 lasing in
  
  31   01 Mar 2025, 06:56 Tokushi SatoSPBShift summaryX-ray delivery

    up to 202/352
pulses per train (~0.8 mJ pulse energy)@13.4
  
  30   28 Feb 2025, 06:57 Tokushi SatoSPBShift summaryX-ray delivery

    up to 202/352
pulses per train (~0.8 mJ pulse energy)@13.4
  
  29   27 Feb 2025, 07:02 Tokushi SatoSPBShift summaryX-ray delivery

    up to 202/352
pulses per train (~0.8 mJ pulse energy)@13.4
  
  28   25 Feb 2025, 02:41 Jayanath KoliyaduSPBShift summaryX-ray delivery
-15 keV with Hirex monitoring, 0.5
mJ/pulse, 202 pulses/train at 0.56 MHz pulse
  
  27   21 Feb 2025, 23:21 Mohammad VakiliSPBShift summarySPB/SFX shift summary user experiment 8039

**X-ray delivery**
-15 keV with Hirex monitoring, 0.5
  
  26   20 Feb 2025, 23:29 Chan KimSPBShift summaryX-ray delivery


12.4 keV, 1.1mJ/pulse,
202 pulses/train at 0.56MHz
Focused
  
  25   19 Feb 2025, 21:25 Chan KimSPBShift summaryX-ray delivery
-12.4 keV, 202 pulses, 0.5 MHz, 1mJ/pulse

Optical laser delivery
  
  24   17 Feb 2025, 06:57 Tokushi SatoSPBShift summaryX-ray delivery

    up to 202 pulses
per train (~1.9 mJ pulse energy)@9.3 keV,
  
  23   16 Feb 2025, 06:46 Tokushi SatoSPBShift summaryX-ray delivery

    up to 202 pulses
per train (~1.9 mJ pulse energy)@9.3 keV,
  
  22   15 Feb 2025, 06:52 Tokushi SatoSPBShift summaryX-ray delivery

    up to 202 pulses
per train (~1.9 mJ pulse energy)@9.3 keV,
  
  21   13 Feb 2025, 07:24 Tokushi SatoSPBShift summaryX-ray delivery

    up to 352 pulses
per train (~1.9 mJ pulse energy)@9.3 keV,
  
  20   11 Feb 2025, 01:41 Tomas PopelarSPBShift summarySPB/SFX 7955 Day 4

X-ray delivery



up to 202 pulses
  
  Draft   10 Feb 2025, 00:56 Raphael de WijnSPBShift summarySPB/SFX 7955 Day
3

X-ray
  
  18   09 Feb 2025, 06:35 Peter ZaldenFXEShift summaryPreparation for user expt #7944

X-ray delivery


Single pulses (10 Hz) at
  
  17   08 Feb 2025, 23:44 Raphael de WijnSPBShift summarySPB/SFX 7955 Day 2

**X-ray delivery**
* up to 202 pulses per train (~1.4
  
  16   08 Feb 2025, 06:25 Peter ZaldenFXEShift summaryPreparation for user expt #7944

X-ray delivery


Since pulses (10 Hz) at
  
  15   08 Feb 2025, 00:00 Tokushi SatoSPBShift summarySPB/SFX 7955 Shift 1
**X-ray delivery**
* up to 202 pulses per train (~1.4
  
  14   07 Feb 2025, 05:56 Peter ZaldenFXEShift summaryPreparation for user expt #7944

X-ray delivery


since pulses (10 Hz) at
  
  13   06 Feb 2025, 23:24 Raphael de WijnSPBIssueDamage on FEL imager was also seen by our
morning shift around 7:50am.
 Screenshot_from_2025-02-06_07-48-20.png 
  12   06 Feb 2025, 23:13 Raphael de WijnSPBShift summaryX-ray delivery


up to 202 pulses per train
(~1.4 mJ pulse energy), 0.5 MHz.


Optical
  
  11   06 Feb 2025, 23:02 Peter ZaldenFXEIssueWe observed a damage spot on the SA1 FEL
imager. It moves with the target holder,
but not sure if its damage on the optical
 2025-02-06-230215_634787810_exflqr51474.png 
  10   05 Feb 2025, 23:42 Raphael de WijnSPBShift summaryX-ray delivery


up to 202 pulses per train
(~1 mJ pulse energy), 0.5 MHz.


Optical
  
  9   01 Feb 2025, 15:40 Frederik Wolff-FabrisPRCStatusSaturday 01.Feb around 2AM machine went
down due to problems with cold magnet at
A20:
  
  8   31 Jan 2025, 21:24 Chan KimSPBShift summaryX-ray delivery


11.0 keV, up to 352 pulses
at 1.1 MHz, 900 uJ.


Optical
  
  7   31 Jan 2025, 12:43 Frederik Wolff-FabrisPRCSA3 statusA ZZ is organized to happen now, 31.01
at ~12:40, in order to reset a cabinet belonging
to one of the Apple-X undulators.
  
  6   31 Jan 2025, 06:28 Jörg HallmannMIDShift summary
Alignment of KB mirrors: focused
beam H150-200nm x V300-350nm. Characterization
of focused beam with wire scan.
Installed
  
  5   30 Jan 2025, 23:08 Chan KimSPBShift summaryX-ray delivery


11.0 keV, up to 178 pulses
at 564 kHz, 650 uJ.


Optical
  
  4   30 Jan 2025, 13:59 Frederik Wolff-FabrisPRCSA3 statusOn 29.01 evening the SA3 pre-absorber was
interlocked due to M5 mirror.

With the help of Maurizio V., it
  
  3   29 Jan 2025, 09:08 Andreas KochXPDSA3 statusThe SA3 FEL imager is now protected via
the Big Brother server, which creates a beam
limit on the bunch pattern server if necessary.
  
  2   19 Jan 2025, 11:46 Jan GruenertPRCIssue1) BKR calls SUN11h20 and informs
about issues with the SA1 SRA. Motors in
ERROR.
 JG_2025-01-20_um_10.24.01.png 
  1   14 Jan 2025, 07:38 Andreas KochXPDSA3 statusThe image intensifier in the SQS Exit Slit
imager had been exchanged on request of SQS,
Thomas Baumann.
  
ELOG V3.1.4-7c3fd00